OpenStudy (anonymous):

What operation is necessary to solve the equation that represents the statement: two point four is 8 multiplied by what number? Explain your answer.

OpenStudy (texaschic101):

write it like you read it... 2.4 is 8 multiplied by what number ok..." is " means equals.....and since we do not know what number, we will use a variable to represent " what number ". we will use n to represent " what number ". so now we have : 2.4 = 8 * n.....same as 2.4 = 8n now we need to solve for n...to get a variable by itself, we do the opposite of what is shown. 8 * n....so multiplication is shown...we do the opposite and divide. divide both sides by 8, which cancels out the 8 on the right side giving us : 2.4/8 = n 0.3 = n so 2.4 = 8 * 0.3